OpenStudy (anonymous):

5 is 2/3 of what number?

OpenStudy (anonymous):

\[\frac{2}{3}\times? =5\] \[?=5\times \frac{3}{4}=\frac{15}{4}\]

OpenStudy (anonymous):

\[\frac{2}{3}*\frac{15}{2}=5 \]